drug usage in the workplace costs employers an incredible amount of money each year. Drug testing potential employees has become so prevalent that drug users are finding it exceptionally difficult to land jobs. Drug tests, however, are not completely reliable. The most common test used to detect the drugs is approximately 99% accurate to decrease the likelihood of making an error, all potential employees are screen through two tests which are independent, and each has about 99% accuracy.


If a person were a drug user, what is the probability that he or she would pass at least one test? (Round 4 places)
Answer #1

P(Pass through atleast one test)

= 1 - P(Fail in both)

= 1 - P(Test accurate for both the cases)

= 1 - 0.99*0.99

= 0.0199